OpenStudy (anonymous):

Find the complete solution for x≡7 (mod 10) and x≡5 (mod 12). Please show steps. @-------College Al…

OpenStudy (anonymous):

This is what I have so far:

OpenStudy (anonymous):

x≡5 (mod 12) -> x= 5 + 12t

OpenStudy (anonymous):

and then I substituted that equation into x≡ 7 (mod 10) to give 5 + 12t = 7(mod 10)

OpenStudy (anonymous):

which gave 12t ≡ 2 (mod 10)

OpenStudy (anonymous):

i'm not sure if that's right, but now i'm stuck. :(

OpenStudy (anonymous):

damn i did it the other way. i said \[x\equiv 7(10)\implies x = 7+10k\] \[7+10k\equiv5(12)\implies 10k\equiv -2(12)\equiv 10(12)\] \[ \implies k \equiv 1(12)\implies k=1+10j\] back substitute to get \[x=7+10k=7+10(1+12j)=7+10+120j=17+120j\]

OpenStudy (anonymous):

i can retry it starting with \[x\equiv 5(12)\] if you like. should work out the same

OpenStudy (anonymous):

that would actually be great, thx :)

OpenStudy (anonymous):

yes it will do what you did exactly. when you get to the step that says \[12t\equiv 2(10)\] then add ten to the left to get \[12t\equiv 12(10)\] so \[t\equiv 1(10)\]

OpenStudy (anonymous):

i meant add ten to the right

OpenStudy (anonymous):

now you have \[t=1+10j\] so \[5+12t=5+12(1+10j)=17+120j\] just like i got

OpenStudy (anonymous):

you can check for a few values of j to see that this is right. for example if \[j=1\] you get 137 which works since it has a remainder of 7 when divided by 10 and a remainder of 5 when divided by 12

OpenStudy (anonymous):

Thank you so much! I've been trying to understand this for weeks! :)

OpenStudy (anonymous):

hope that helped. i think the key step might have been starting with \[12t\equiv 2(10)\] and then getting \[112t\equiv 1(10)\]

OpenStudy (anonymous):

yw

OpenStudy (stacey):

x=17+60n If you use 120, where I have 60, you will miss solutions such as x=77.

OpenStudy (anonymous):

Thanks, Stacey! I just checked it over.

OpenStudy (stacey):

The error occurs with 12t=12 (mod 10) When you divide by 12 the modulus is divided by any common factor. t=1 (mod 5)